# The $\infty$-S test via quantile affine LASSO

### v2

@@ 3 Optimization
The proof is straightforward: choose $m$ linearly independent columns of $A$, call them $A_{1}$ and the remaining ones $A_{2}$, so that $A=[A_{1},A_{2}]$, permuting $\boldsymbol{\beta}=(\boldsymbol{\beta}_{1},\boldsymbol{\beta}_{2})$ accordingly. Hence the linear constraints leads to $\boldsymbol{\beta}_{1}=A_{1}^{-1}(\boldsymbol{b}-A_{2}\boldsymbol{\beta}_{2})$ and $\boldsymbol{y}-X\boldsymbol{\beta}$ becomes $\boldsymbol{y}_{-A}-X_{-A}\boldsymbol{\beta}_{2}$ with $\boldsymbol{y}_{-A}=\boldsymbol{y}-X_{1}A_{1}^{-1}\boldsymbol{b}$ and $X_{-A}=X_{2}-X_{1}A_{1}^{-1}A_{2}$.
The affine LAD-LASSO optimization (5) can also be transformed into an unpenalized LAD of the form (3) since it is equivalent to solving
$\min_{{\boldsymbol{\beta}}\in{\mathbb{R}}^{p}}\|\left(\begin{array}[]{c}{\boldsymbol{y}}\\ {\lambda{\boldsymbol{b}}}\end{array}\right)-\left(\begin{array}[]{c}X\\ {\lambda A}\end{array}\right){\boldsymbol{\beta}}\|_{1}.$
Another way to solve the constrained LAD optimization (6) is to solve the affine LAD LASSO (5) for $\lambda=\lambda_{0}(\boldsymbol{Y},X,A)$ of Theorem 1 to guarantee complete thresholding, that is, forcing the linear constraint; this is akin to the exact penalty method of Di Pillo and Grippo (1989). The following theorem links the dual of the affine LAD LASSO optimization (5) and the constrained LAD optimization.
@@
The dual variable $\boldsymbol{\omega}$ of Theorem 1 corresponding to the primal affine LAD LASSO optimization (5) is the dual variable of the LAD optimization (8) with $\boldsymbol{y}_{-A}$ and $X_{-A}$ of Theorem 3.
So the $\infty$-S test can straightforwardly be implemented with current softwares, in particular using the quantreg package (Koenker, 2024) in R.
4 The quantile $\infty$-Sτ test
Given $\tau\in(0,1)$, consider the $\tau$-conditional quantile vector ${\boldsymbol{q}}^{\tau}$ of a response vector ${\boldsymbol{Y}}\mid X$, where $X$ is the regression matrix. The linear model assumes
${\boldsymbol{q}}^{\tau}=X{\boldsymbol{\beta}}.$ (9)
To estimate the parameters ${\boldsymbol{\beta}}$ based on a sample ${\boldsymbol{y}}$ from ${\boldsymbol{Y}}\mid X$, quantile regression (Koenker and Bassett, 1978) considers solving
$\min_{\boldsymbol{\beta}\in{\mathbb{R}}^{p}}\|\boldsymbol{y}-X\boldsymbol{\beta}\|_{\rho_{\tau}},$ (10)
where the residuals are $\boldsymbol{r}=\boldsymbol{y}-X\boldsymbol{\beta}$ and $\|\boldsymbol{r}\|_{\rho_{\tau}}=\sum_{i=1}^{n}\rho_{\tau}(r_{i})$ with
$\rho_{\tau}(r_{i})=\left\{\begin{array}[]{rl}(\tau-1)r_{i}&r_{i}<0\\ \tau r_{i}&r_{i}\geq 0\end{array}\right.$ (11)
is the tilted $\ell_{1}$ loss. In particular $\|\boldsymbol{r}\|_{\rho_{1/2}}=\|{\boldsymbol{r}}\|_{1}/2$.
To test the null hypothesis against the alternative hypothesis (2), in the spirit of quantile regression, one derives the quantile $\infty$-S test to not necessarily fit the data to the median (that is $\tau=1/2$), but to any $\tau$-quantile for some fixed $\tau\in(0,1)$. To that aim, we define affine $\rho_{\tau}$-LASSO as a solution to
$\min_{\boldsymbol{\beta}\in{\mathbb{R}}^{p}}\rho_{\tau}(\boldsymbol{y}-X\boldsymbol{\beta})+\lambda\|A\boldsymbol{\beta}-\boldsymbol{b}\|_{1}$ (12)
for some $\lambda>0$ and $\tau\in(0,1)$. This estimator has a zero-thresholding function given by the following theorem.
@@ 4 The quantile $\infty$-Sτ test
Given $\tau\in(0,1)$, let
$\hat{\boldsymbol{\beta}}_{H_{0}}^{\tau}\in\arg\min_{\boldsymbol{\beta}\in{\mathbb{R}}^{p}}\rho_{\tau}(\boldsymbol{y}-X\boldsymbol{\beta})\quad{\rm s.t.}\quad A\boldsymbol{\beta}=\boldsymbol{b}.$ (13)
@@
where ${\boldsymbol{\omega}}^{\tau}$ is the dual variable associated to (12), which are the weighted sign of $X\hat{\boldsymbol{\beta}}_{H_{0}}^{\tau}-{\boldsymbol{y}}$ of the constrained minimum $\rho_{\tau}$ (13).
The zero-thresholding function leads to the statistic of the $\infty$-Sτ test, which is asymptotically pivotal, as stated in the following theorem.
@@
Given $\tau\in(0,1)$, assuming the quantile regression coefficient estimates are asymptotically Gaussian centered around the true coefficients, the test statistic $S^{\tau}=\lambda_{0}^{\tau}(\boldsymbol{Y},X,A)$, where $\boldsymbol{Y}$ is the response random vector under $H_{0}$ is asymptotically pivotal.
Asymptotic Gaussianity has been proved by Koenker and Bassett (1978), so the $\infty$-Sτ test can now be defined.
@@
Given $\tau\in(0,1)$, consider the linear model (9) for the quantiles. To test (2) at a prescribed level $\alpha\in(0,1)$, let $c_{\alpha}^{\tau}=F_{S^{\tau}}^{-1}(1-\alpha)$, where $S^{\tau}=\lambda_{0}^{\tau}(\boldsymbol{Y},X,A)$ and $\boldsymbol{Y}$ is the response random vector for the linear model under $H_{0}$ and the function $\lambda_{0}^{\tau}$ is given in (14). The quantile $\infty$-Sτ test function is defined by
$\phi({\bf y},X,A)=\left\{\begin{array}[]{rl}0&\lambda_{0}^{\tau}(\boldsymbol{y},X,A)\leq c_{\alpha}^{\tau}\\ 1&{\rm otherwise}\end{array}\right.,$
where ${\boldsymbol{y}}$ is the data response.
Based on Theorem 6, the level of the $\infty$-Sτ test is asymptotically the nominal level $\alpha$. Weighting the sign $\boldsymbol{\omega}^{\tau}$ of the residuals by the ranks of the weighted residuals $|\boldsymbol{r}|\cdot\rho_{\tau}(\boldsymbol{r})$ leads to the $\infty$-rankSτ test.
5 Application
Our simulations and application are done with the R software and rely on the quantreg package (Koenker, 2024). We use the rq function to calculate the dual variables $\boldsymbol{\omega}$ necessary to calculate the zero-thresholding function. The critical values are estimated by Monte Carlo simulation with $10^{4}$ runs in all our applications of the $\infty$-S test. The levels and power curves of the tests considered are estimated based on $10^{4}$ simulated data sets. We use the anova.rq function to perform the $\chi^{2}$ sign test whenever it returned a p-value; otherwise the results of the $\chi^{2}$ sign test cannot be reported.
@@ 5.1 The paired $\infty$-S test
with the $\infty$-S test, one writes the model as ${\boldsymbol{y}}=X\beta+{\boldsymbol{\epsilon}}$ with ${\boldsymbol{y}}={\boldsymbol{v}}-{\boldsymbol{u}}$, $X={\boldsymbol{1}}_{n}$, $\beta=\delta$, ${\boldsymbol{\epsilon}}={\tilde{\boldsymbol{\epsilon}}}-{\breve{\boldsymbol{\epsilon}}}$, $A=1$ and $b=0$. Then the test-statistic value is $s=|\sum_{i=1}^{n}{\rm sign}({\boldsymbol{y}})|=|\sum_{i=1}^{n}{\rm sign}({\boldsymbol{v}}-{\boldsymbol{u}})|=|\sum_{i=1}^{n}(1(v_{i}>u_{i})-1(v_{i}<u_{i}))|$. Recall that the sign test’s statistic value is $\tilde{s}=\sum_{i=1}^{n}1(v_{i}>u_{i})$. But $\sum_{i=1}^{n}1(v_{i}>u_{i})+\sum_{i=1}^{n}1(v_{i}<u_{i})=n$, so $s=|2\tilde{s}-n|$ and the $\infty$-S test is equivalent to the existing sign test.
@@ 5.2 The unpaired $\infty$-S test
${\boldsymbol{\beta}}=(\mu,\delta)^{\rm T}$, ${\boldsymbol{\epsilon}}^{\rm T}=({\tilde{\boldsymbol{\epsilon}}}^{\rm T},{\breve{\boldsymbol{\epsilon}}}^{\rm T})$, $A=(0,1)$ and $b=0$. One gets $\hat{\boldsymbol{\beta}}_{H_{0}}=(m_{\boldsymbol{y}},0)^{\rm T}$, and the test-statistic value is $s=\max(0,|\sum_{i=1}^{n}1(v_{i}>m_{\boldsymbol{y}})-\sum_{i=1}^{n}1(v_{i}<m_{\boldsymbol{y}})|)$. This is the sign test for unpaired data, also known as the median test (Brown and Mood, 1951).
@@ 5.3 Homopower rescaling
We perform a toy simulation to outline the importance of homopower rescaling (see Definition 2) the two components $A$ and $\boldsymbol{b}$ of the linear null hypothesis (2). Here $n=100$, $p=20$, $m=2$, $\tau=0.5$, $\boldsymbol{b}=\boldsymbol{0}_{m}$ and $A=[C,O_{m\times(p-m)}]$ with
@@
The regression matrix $X$ has entries sampled from a standard Gaussian random variable, the null model is $H_{0}:\ \boldsymbol{\beta}=\boldsymbol{0}$, and we consider two alternative hypotheses
@@
where $\boldsymbol{e}_{i}$ is the $i$th canonical vector. With the unrescaled $\infty$-S test, one expects the power to be low under $H_{1,1}$ and good under $H_{1,2}$, while with the rescaled $\infty$-S test, one expects the power to be high under both $H_{1,1}$ and $H_{1,2}$. Figure 2 illustrates the improvement thanks to homopower rescaling. Not only the $\infty$-S test, but many tests would profit from this rescaling in terms of power. This example is certainly a toy example, but what we observe here also occurs naturally owing to the relation between $A$ and $X$ that is not controlled by the user, except possibly in ANOVA. So homopower rescaling should be employed.
@@ 5.4 The total variation $\infty$-Sτtest
Total variation (Rudin et al., 1992) aims at detecting jumps in a noisy signal or a time series ${\bf y}=\boldsymbol{\beta}+\boldsymbol{e}$, which corresponds to model (1) with $X=I_{n}$. A jump occurs if $\sum_{j=1}^{n-1}|\beta_{j+1}-\beta_{j}|\neq 0$. Call $A$ the $(n-1)\times n$ sparse matrix with $-1$ and $1$ entries such that $\sum_{j=1}^{n-1}|\beta_{j+1}-\beta_{j}|\ =\|A\boldsymbol{\beta}\|_{1}$; $A$ is full row rank. To test for jump in the median or any $\tau$-quantile, one can apply the $\infty$-Sτ test or the $\infty$-rankSτ test at the desired $\tau$-quantile.
We apply the quantile total variation $\infty$-Sτ test to analyse the time series of the amazon daily log-returns. Focusing on the lower tail which corresponds to drops in the AMZN stock value, we test for constant $\tau$-quantile for $\tau=0.1$. The $\infty$-Sτ test is rejected with an estimated p-value of $0.0132$. The top left plot of Figure 3 shows the histogram estimate of the density $f_{S^{\tau}}$ of the test statistic (by Monte Carlo sampling of the test statistic $S^{\tau}$ under the null hypothesis) and the critical value $c_{\alpha}^{\tau}$ for $\alpha=0.05$. Right below this plot, Figure 3 shows the quantile affine LASSO path, along with the critical value $c_{\alpha}^{\tau}$ and the zero-thresholding function $\lambda_{0}^{\tau}(\boldsymbol{y},X,A)$ evaluated at the data $\boldsymbol{y}$. The right plots of Figure 3 show the raw AMZN time series data (top) and the quantile affine LASSO estimate for $\lambda=c_{\alpha}^{\tau}$ (bottom), which corresponds to the quantile universal threshold estimate (Giacoetal17) of the $\tau$-quantile. The single jump points to a potential change of regime moving to a less volatile return.
Figure 3: AMZN daily log-returm time series analysis. Top left plot: estimated density function of $f_{S^{\tau}}$ and critical value $c_{\alpha}^{\tau}$ for $\alpha=0.05$ by Monte Carlo simulation. Bottom left: regression quantile affine LASSO path for $\tau=0.1$ with critical value $c_{\alpha}^{\tau}$ and zero-thresholding function $\lambda_{0}(\boldsymbol{y},X,A)$ evaluated at $\boldsymbol{y}$. Top right plot: AMZN time series data and estimate jump locations. Bottom right: estimated $\tau$-quantiles.

### v3

@@ 3 Optimization aspects
The proof is straightforward: choose $m$ linearly independent columns of $A$, call them $A_{1}$ and the remaining ones $A_{2}$, so that $A=[A_{1},A_{2}]$, permuting $\boldsymbol{\beta}=(\boldsymbol{\beta}_{1},\boldsymbol{\beta}_{2})$ accordingly. The linear constraints thus lead to $\boldsymbol{\beta}_{1}=A_{1}^{-1}(\boldsymbol{b}-A_{2}\boldsymbol{\beta}_{2})$ and $\boldsymbol{y}-X\boldsymbol{\beta}$ becomes $\boldsymbol{y}_{-A}-X_{-A}\boldsymbol{\beta}_{2}$ with $\boldsymbol{y}_{-A}=\boldsymbol{y}-X_{1}A_{1}^{-1}\boldsymbol{b}$ and $X_{-A}=X_{2}-X_{1}A_{1}^{-1}A_{2}$.
The affine LAD-LASSO optimization (4) can be also transformed into an unpenalized LAD of the form (3), as it is equivalent to solving
$\min_{{\boldsymbol{\beta}}\in{\mathbb{R}}^{p}}\biggl\|\left(\begin{array}[]{c}{\boldsymbol{y}}\\ {\lambda{\boldsymbol{b}}}\end{array}\right)-\left(\begin{array}[]{c}X\\ {\lambda A}\end{array}\right){\boldsymbol{\beta}}\biggr\|_{1}.$
Another way to perform the constrained LAD optimization (5) is to solve the affine LAD-LASSO (4) for $\lambda=\lambda_{0}(\boldsymbol{Y},X,A)$ of Theorem 1 to guarantee complete thresholding—that is, enforcing the linear constraint; this is akin to the exact penalty method of PG89. The following theorem links the dual of the affine LAD-LASSO optimization (4) and the constrained LAD optimization.
@@
The dual variable $\boldsymbol{\omega}$ of Theorem 1 corresponding to the primal affine LAD LASSO optimization (4) is the dual variable of the LAD optimization with $\boldsymbol{y}_{-A}$ and $X_{-A}$ of Theorem 3.
Summing up, the $\infty$-S test can be straightforwardly implemented using the existing software—in particular, employing the R package quantreg (Rogerrq).
4 The $\infty$-Sτ test for quantile regression
Given $\tau\in(0,1)$, let ${\boldsymbol{q}}^{\tau}$ be conditional $\tau$-quantile vector ${\boldsymbol{q}}^{\tau}$ of a response vector ${\boldsymbol{Y}}\mid X$. The linear model assumes the form
${\boldsymbol{q}}^{\tau}=X{\boldsymbol{\beta}},$ (8)
where $X$ is the regression matrix as before. To estimate the parameters $\boldsymbol{\beta}$} from $\boldsymbol{y}$} arising from the response random vector $\boldsymbol{Y}$}, quantile regression (KB78) solves the optimization problem
$\|\boldsymbol{y}-X\boldsymbol{\beta}\|_{\rho_{\tau}}\to\min_{\boldsymbol{\beta}\in{\mathbb{R}}^{p}}!$ (9)
The residuals $\boldsymbol{r}=\boldsymbol{y}-X\boldsymbol{\beta}$ are now subjected to the objective function $\|\boldsymbol{r}\|_{\rho_{\tau}}=\sum_{i=1}^{n}\rho_{\tau}(r_{i})$, where
$\rho_{\tau}(r_{i})=\left\{\begin{array}[]{rl}(\tau-1)r_{i}&\ r_{i}<0,\\ \tau r_{i}&\ r_{i}\geq 0,\end{array}\right$. (10)
is the “check function”—the tilted $\ell_{1}$, “pinball” loss. Note that the special case of $\tau=1/2$ reduces to the LAD regression, as $\|\boldsymbol{r}\|_{\rho_{1/2}}=\|{\boldsymbol{r}}\|_{1}/2$.
To test the null hypothesis against the alternative hypothesis, as in (2) but in the quantile-regression spirit, the quantile $\infty$-S test does not refer to the median (well, unless $\tau=1/2$), but rather to (any) $\tau$-quantile for some fixed $\tau\in(0,1)$. We define accordingly the affine $\rho_{\tau}$-LASSO as a solution to
$\rho_{\tau}(\boldsymbol{y}-X\boldsymbol{\beta})+\lambda\|A\boldsymbol{\beta}-\boldsymbol{b}\|_{1}\to\min_{\boldsymbol{\beta}\in{\mathbb{R}}^{p}}!$ (11)
for fixed $\lambda>0$ and $\tau\in(0,1)$. The following theorem gives the zero-thresholding function of this estimator.
@@ 4 The $\infty$-Sτ test for quantile regression
For fixed $\tau\in(0,1)$, let
$\hat{\boldsymbol{\beta}}_{H_{0}}^{\tau}\in\arg\min_{\boldsymbol{\beta}\in{\mathbb{R}}^{p}}\rho_{\tau}(\boldsymbol{y}-X\boldsymbol{\beta})\quad{\rm subject\ to}\quad A\boldsymbol{\beta}=\boldsymbol{b}.$ (12)
@@
where ${\boldsymbol{\omega}}^{\tau}$ is the dual variable associated to the linear programming solution of (11).
The zero-thresholding function leads to the statistic of the $\infty$-Sτ test; the following theorem shows that it is asymptotically pivotal.
@@
Let $\boldsymbol{Y}$ be the response random vector under $H_{0}$ and let $\tau\in(0,1)$. If the $\tau$-quantile regression parameter estimates are asymptotically normal with the limit distribution centered about their true values, the test statistic $S^{\tau}=\lambda_{0}^{\tau}(\boldsymbol{Y},X,A)$ is asymptotically pivotal.
The asymptotic normality of the quantile regression estimators, as established by KB78 (their results including the $\ell_{1}$ results as a special case), allows for the following definition of the $\infty$-Sτ test—which Theorem 6 shows has asymptotically the nominal level $\alpha$.
@@
Given $\tau\in(0,1)$, the $\infty$-Sτ test to test (2) in the linear model (8) at a prescribed level $\alpha\in(0,1)$, is defined to be
$\phi({\boldsymbol{y}},X,A)=\left\{\begin{array}[]{rl}0&\lambda_{0}^{\tau}(\boldsymbol{y},X,A)\leq c_{\alpha}^{\tau}\\ 1&{\rm otherwise}\end{array}\right$.,
where $c^{\tau}_{\alpha}$ is selected so that the test has nominal level $\alpha$.
5 Special cases and applications
Our simulations and applications rely on the R package quantreg (Rogerrq). We use the rq function to calculate the dual variables $\boldsymbol{\omega}$ involved in the zero-thresholding function. In all our applications of the $\infty$-S test, the critical values are estimated by Monte Carlo simulation with $10^{4}$ runs. The levels and power curves of the tests are estimated based on $10^{4}$ simulated data sets. We use the anova.rq function to perform the $\chi^{2}$ sign test whenever it returns a p-value; otherwise the results of the $\chi^{2}$ sign test cannot be reported.
@@ 5.1 The paired $\infty$-S test
with the $\infty$-S test, we write the model as ${\boldsymbol{y}}=X\beta+{\boldsymbol{\epsilon}}$ with ${\boldsymbol{y}}={\boldsymbol{v}}-{\boldsymbol{u}}$, $X={\boldsymbol{1}}_{n}$, $\beta=\delta$, ${\boldsymbol{\epsilon}}={\tilde{\boldsymbol{\epsilon}}}-{\breve{\boldsymbol{\epsilon}}}$, $A=1$ and $b=0$. The value of the test statistic is then
$s=|\sum_{i=1}^{n}{\rm sign}({\boldsymbol{y}})|=|\sum_{i=1}^{n}{\rm sign}({\boldsymbol{v}}-{\boldsymbol{u}})|=|\sum_{i=1}^{n}(1(v_{i}>u_{i})-1(v_{i}<u_{i}))|.$
Given that the statistic of the sign test statistic in this situation is
$\tilde{s}=\sum_{i=1}^{n}1(v_{i}>u_{i})\quad\text{and}\quad\sum_{i=1}^{n}1(v_{i}>u_{i})+\sum_{i=1}^{n}1(v_{i}<u_{i})=n,$
we have that $s=|2\tilde{s}-n|$ and the $\infty$-S test is equivalent to the sign test.
@@ 5.2 The unpaired $\infty$-S test
${\boldsymbol{\beta}}=(\mu,\delta)^{\rm T}$, ${\boldsymbol{\epsilon}}^{\rm T}=({\tilde{\boldsymbol{\epsilon}}}^{\rm T},{\breve{\boldsymbol{\epsilon}}}^{\rm T})$, $A=(0,1)$ and $b=0$. One gets now $\hat{\boldsymbol{\beta}}_{H_{0}}=(m_{\boldsymbol{y}},0)^{\rm T}$, and the test-statistic value is
$s=\max(0,|\sum_{i=1}^{n}1(v_{i}>m_{\boldsymbol{y}})-\sum_{i=1}^{n}1(v_{i}<m_{\boldsymbol{y}})|).$
We obtain the sign test for unpaired data, also known as the median test (brown1951median).
@@ 5.3 Homopower rescaling
We perform a small simulation study to outline the importance of homopower rescaling (see Definition 2) of the two components $A$ and $\boldsymbol{b}$ of the linear null hypothesis (2). We set $n=100$, $p=20$, $m=2$, $\tau=0.5$, $\boldsymbol{b}=\boldsymbol{0}_{m}$ and $A=[C,O_{m\times(p-m)}]$ with
@@
The regression matrix $X$ has entries sampled from a standard normal random distribution, the null hypothesis is $H_{0}:\ \boldsymbol{\beta}=\boldsymbol{0}$; we consider two alternative hypotheses
@@
where $\boldsymbol{e}_{i}$ is the $i$th coordinate (canonical) vector. With the unrescaled $\infty$-S test, one expects the power to be low under $H_{1,1}$ and good under $H_{1,2}$, while with the rescaled $\infty$-S test, one expects the power to be high under both $H_{1,1}$ and $H_{1,2}$. Figure 2 illustrates that it is really so. While this example represents certainly just one very special case, we remark that what we observe here also occurs naturally, owing to the relation between $A$ and $X$ that is not controlled by the user (except possibly in ANOVA). Thus, it is not only the $\infty$-S test, but many other tests that would profit from such a rescaling in terms of power.
@@ 5.4 The total variation $\infty$-Sτtest
Total variation (ROF92) aims at detecting jumps in a noisy signal or a time series ${\bf y}=\boldsymbol{\beta}+\boldsymbol{e}$, which corresponds to model (1) with $X=I_{n}$. A jump occurs if $\sum_{j=1}^{n-1}|\beta_{j+1}-\beta_{j}|\neq 0$. Let $A$ be the $(n-1)\times n$ (sparse) matrix with 1-1 and $1$ entries such that
$\sum_{j=1}^{n-1}|\beta_{j+1}-\beta_{j}|\ =\|A\boldsymbol{\beta}\|_{1}.$
Note that $A$ is of full row rank. To test for jump in the median or any $\tau$-quantile, one can apply the $\infty$-Sτ test or the $\infty$-rankSτ test at the desired $\tau$-quantile.
We apply the quantile total variation $\infty$-Sτ test to analyze the time series of the Amazon daily log-returns. Focusing on the lower tail which corresponds to drops in the AMZN stock value, we test for constant $\tau$-quantile for $\tau=0.1$. The $\infty$-Sτ test is rejected with an estimated p-value of $0.0132$. The top left plot of Figure 3 shows the histogram estimate of the density $f_{S^{\tau}}$ of the test statistic (by Monte Carlo sampling of the test statistic $S^{\tau}$ under the null hypothesis) and the critical value $c_{\alpha}^{\tau}$ for $\alpha=0.05$. Right below this plot, Figure 3 shows the quantile affine LASSO path, along with the critical value $c_{\alpha}^{\tau}$ evaluated at the data $\boldsymbol{y}$. The right plots of Figure 3 show the raw AMZN time series data (top) and the quantile affine LASSO estimate for $\lambda=c_{\alpha}^{\tau}$ (bottom), which corresponds to the quantile universal threshold estimate (Giacoetal17) of the $\tau$-quantile. The single jump points to a potential change of regime moving to a less volatile return.
Figure 3: AMZN daily log-returm time series analysis. Top left plot: estimated density function of $f_{S^{\tau}}$ and critical value $c_{\alpha}^{\tau}$ for $\alpha=0.05$ by Monte Carlo simulation. Bottom left: regression quantile affine LASSO path for $\tau=0.1$ with critical value $c_{\alpha}^{\tau}$. Top right plot: AMZN time series data and estimate jump locations. Bottom right: estimated $\tau$-quantiles.